OpenStudy (anonymous):

(sin168 - sin108)/(cos168+cos108)

OpenStudy (anonymous):

2.cos(168+108/2).sin(30) / 2.cos(168+108/2).cos(30) sin(30)/cos(30)= tan(30)= 1/[\sqrt{3}\]

OpenStudy (anonymous):

1/sqrt(3)